OpenStudy (anonymous):

What is c * f(x), if f(x)= x^2 and c= 5?

OpenStudy (anonymous):

I know that c * f(x) indicates that the entire existing function will be multiplied by a constant but I am not sure how the function would look like

OpenStudy (anonymous):

For example: f(cx), if f(x) = x^2 and c = 5, then f(cx) = (5x)^2. Also what is the difference between f(cx) and c * f(x)?

OpenStudy (anonymous):

@Hero do you think you can help me once again? I'm sorry to bother you...

OpenStudy (anonymous):

@ash2326 or @mayankdevnani I am sorry, it is that I really need help

OpenStudy (mayankdevnani):

you mean that \[\huge \bf c \times f(x)\]

OpenStudy (ash2326):

@issy987 Are you here?

OpenStudy (anonymous):

Yes, sorry my computer crashed but I am here

OpenStudy (ash2326):

you asked how will a function look like if it's multiplied by a constant?

OpenStudy (mayankdevnani):

@issy987 is that your question what i have written?

OpenStudy (anonymous):

Yes, I am looking for c * f(x) when f(x) is x^2 and c is 5

OpenStudy (mayankdevnani):

just multiply them

OpenStudy (mayankdevnani):

just do simple multiplication

OpenStudy (mayankdevnani):

okk.. @issy987

OpenStudy (anonymous):

But wouldn't it become (5x)^2, because f(cx) = (5x)^2

OpenStudy (mayankdevnani):

you are wrong. its 5x^2

OpenStudy (mayankdevnani):

where, 5=c and f(x)=x^2

OpenStudy (mayankdevnani):

i hope you understand.

OpenStudy (anonymous):

Ya, for f(cx) the answer was in the parentheses and for c * f(x) the answer is not.

OpenStudy (mayankdevnani):

what is the answer by the way?

OpenStudy (anonymous):

for c * f(x) it would be 5x^2 right?

OpenStudy (mayankdevnani):

right

OpenStudy (anonymous):

ok, though I have another question, do you mind if I ask you?

OpenStudy (mayankdevnani):

oh god, of course no